force on femur

The normal force of the ground on the foot can reach three times a runner's body weight when the foot strikes the pavement.

By what amount does the 52 cm long femur of an 67 kg runner compress at this moment? The cross-section area of the bone of the femur can be taken as 5.2X10^-4 m^2.
Answer #1

Quite simple actually   


F=Y*A/L*ΔL

We can rewrite it like this.


ΔL=F*L/Y/A


F=67kg*9.8m/s2

L=.52m

A=5.2*10-4

Y=1.6*1010

Since the force in multiplied by 3x when the foot hits the ground we multiply the force x3.


F=67kg*9.8m/s2*3


ΔL=(67kg*9.8m/s2*3)*.52/(1.6*1010)/(5.2*10-4)=0.00012 m

change to cm = 0.012 cm

source: Brain
answered by: Joe